OpenStudy (anonymous):

A student pushes horizontally eastward on a 1.2 kg textbook initially resting 1.3 m from the edge of a long table. The coefficient of static friction is 0.24 and the coefficient of kinetic friction is 0.18. The force is just large enough to get the book to start moving, and is maintained for 1.0s after the motion has begun. determine the force applied to the book by the student and determine the acceleration of the book while the student is applying the push.

OpenStudy (anonymous):

Start by getting the weight of the object, which is the force normal to the table, W = m g, where m = 1.2 kg and g = 9.8 N/kg Fs=static friction force = (0.24) (W) Fk = kinetic friction force = (0.18)(W) Pushing the book will require F = Fk + (mass )(acceleration, and f the same force is applied as just = static friction force, then Fs = Fk + m a from which you can get a (acceleration). You might check that the book does not fall off the table, as distance will be d = (1/2) a t^2.